A baseball team has 9 players, all of whom have equal skill in all positions. The team needs 3 outfielders. Assuming order is important, find the number of ways of forming an outfield with 3 outfielders?
Answer by stanbon(75887)

9P3 = 9!/(9-3)! = 9*8*7 = 504
============
Cheers,
Stan H.